OpenStudy (anonymous):

Suppose I and J are ideals in a ring R and let f: R -> R/I X R/J be the function defined by f(a)=(a+I, a+J). a) Prove that f is a homomorphism of rings b) Is f surjective? (Look at the care when R=Z, I=(2) and J=(4)) c) What is the kernel of f?

OpenStudy (anonymous):

write down exactly that you have to prove, that \[f(a+b)=f(a)+f(b)\] which means that \[(a+b)+I, (a+b)+J=((a+I)+(b+I),(a+J) +(b+J))\] which it does, by definition of \((a+I)+(b+I)\)

OpenStudy (anonymous):

left hand side of the above equal sign should be an ordered pair, that is \[\left((a+b)+I,(a+b)+J\right)\]

OpenStudy (anonymous):

then you have to show that \[f(ab)=f(a)f(b)\] which means that \[(ab+I, ab+J)=(a+I, a+J)\times (b+I,b+J)\] which again is true by the definition of multiplication of factor rings

OpenStudy (anonymous):

kernel is everything that gets sent to zero in \(R/I\) the zero element is \(I\) so kernel should be \(K=\{a\in R|a\in I, a\in J\}\) which is another way of saying \(K=I\cap J\)

OpenStudy (anonymous):

\(\mathbb{Z}/(2)=\mathbb{Z}_2=\{0,1\}\) the other one is \(\{0,1,2,3\}\)

OpenStudy (anonymous):

Thank you! I think I have it figured out now!

OpenStudy (anonymous):

yw